A financial company has determined that the price-to-earnings ratios for 20 randomly selected companies ranges between 0.9 and 2.9. Given that the price-to-earnings ratios are uniformly distributed, what proportion (or per cent) of the price-to-earnings ratios are between 1.90 and 2.48?
Answer by jrfrunner(365)

in a uniform probabilty distribution ranging (a, b)
The probability of x1<= x <=x2 = (x2-x1)/(b-a)
---
in your case U(0.9,2.9)
P(1.90<= x <= 2.48)=(2.48-1.90)/(2.9-0.9)=0.58/2=0.49